On January 1 of the current year, Rhondell Corporation has accumulated E & P of $180,000. Current E & P for the year is $540,000, earned evenly throughout the year. Elizabeth and Jonathan are sole equal shareholders of Rhondell from January 1 to April 30. On May 1, Elizabeth sells all of her stock to Marshall. Rhondell makes two distributions to shareholders during the year: a total of $288,000 ($144,000 to Elizabeth and $144,000 to Jonathan) on April 30 and a total of $504,000 ($252,000 to Jonathan and $252,000 Marshall) on December 31.

Determine the allocation of the distributions by completing the table below. Assume that the shareholders have sufficient basis in their stock for any amount that is treated as return of capital.

If an amount is zero, enter "0". If required, round any division to two decimal places and use in subsequent computations. Round final answers to the nearest dollar.

April 30 distribution of $288,000

From Current E & P ______

From Accumulated E & P______

Treated as Return of Capital _________

December 31 distribution of $504,000

From Current E&P______

From Accumulated E&P_______

Treated As Return of Capital_______
Answer #1
April 30 distribution of $288,000
From Current E & P--- 180000--------(540000/12*4)
From Accumulated E & P ----288000-180000= 108000
Treated as Return of Capital------- 0 (Zero)
December 31 distribution of $504,000
From Current E&P ----- 360000
From Accumulated E&P----72000
Treated As Return of Capital ---- 72000------- (504000-360000-72000)
Alternately for Dec 31, distribution:
Dec 31 distribution:
Amt. paid out of total E&P =Total distribution*(Total E&P/Total distribution)
ie. 504000*((360000+72000)/504000)=
432000
Balance --return of capital
ie. 504000-432000=
72000